Director's Time / Engineering

Observers: B.Rothberg, D.Thompson (Tucson)

ISA: B.Rothberg
Telescope Operator: G.Benchetti (LBTO)
AO Support: JCGuerra
Telescope Support: M.Wagner (LBTO)
Software Support: TGolota
Instrument Support: E.Shield (LBTO)
Engineering Support: DAshby

Summary

Details

UT 01:28 - Science integration started J220710

UT 01:46 - Seeing ~0.91 on GCS

UT 02:09 - Seeing 0.86" GCS, 0.81" DIMM

UT 02:25 - Completed J220710

UT 02:26 - Moving to Feige 110 SpectroStd (for 0.6 dual and 5" red only)

UT 02:33 - Feige 110

UT 02:43 - Seeing 0.83" GCS, 1.13" DIMM

UT 02:58 - Moving to PS0J3338p29

UT 03:07 - Science observations started

UT 04:05 - PSC0J3338p29 finished

UT 04:06 - Switching to LUCI1 on SX

UT 04:33 - Moving to SDSSJ0122+0100

UT 04:38 - 0.7" Guider, 0.9" DIMM

UT 04:49 - Acquired, but SX mirror ripped

UT 04:56 - Secondary not settling

UT 05:37 - Reconfigured for a thru-slit image to make sure acq. alignment held

UT 05:40 - Not clear object was still in slit - sending acquisition again

UT 05:55 - Centered in slit - science script started

UT 05:58 - Seeing 0.67" Guider, 0.83" DIMM

UT 06:19 - Not seeing 2nd galaxy spectrum - shift should be +8.5" - aborting on Position 5 in script (centered on left galaxy) - manually shifting +8.5" and taking thru-slit on 2nd galaxy

-current position is -8.5,+30 (image 23)

-now at +8.5,+30 (image 24) - will see if galaxy is back in slit - confirmed, galaxy is back in slit

-now moving -8.5 again to see if this is a mirror in X problem (image 25) - can see edge of right galaxy - looks like part is cut off

-trying +8.2 now to see if more of right galaxy is in slit (image 27)

-going to reacquire - using +/-7.85" for main part of galaxy on right, 8.35 gets blob up top (_fixed.spec) - NOTE: first two images on "right" galaxy is fainter blobs

UT 06:50 - now need to do a pointing check

UT 06:55 - Re-acquiring, with new split script

UT 07:06 - Restarted new script - seeing 0.78" GCS - first science exposure is #32

UT 07:40 - Seeing 0.87" guider

UT 07:53 - stopped at exposure 8 - going to center on galaxy at the right and dither along slit (Now at airmass 2)

UT 08:03 - staring _galright script, - at high airmass - will go for a pair - shell appears to have ripped

UT 08:11 - recovered, guider reads 1.1" - woohoo! first spectra centered on right galaxy and can see broad Pa Alpha - hit elevation limit after pair - can see broad Pa Alpha and Brackett Gamma in data

UT 08:18 - slewing to zenith - recover shell

UT 08:24 - acquiring Telluric std HIP 13917

UT 08:32 - taking Telluric spectra

UT 08:39 - Luci Rothberg Program completed

UT 08:59 - Putting MODS1 to sleep (we will need to get calibrations for Dave's Program - Red Only)

UT 09:11 - in Binocular mode with LUCI1 and MODS2

UT 09:18 - Now testing LUCI1 with SDSSJ0833+1204 - Bulgeless AGN program - sending LUCI1 acquisition while MODS2 deals with pointing

UT 09:20 - LUC1 acq script set up - paused for MODS after "Instrument Setup Successfully Committed" - awaiting MODS2 preset - had to resend MODS2 preset

UT 09:29 - Acquiring LUC1 SDSSJ0833+2040 - need to increase acquisition image time, doing 60x2 and moved to -5,-5 for offset (IMPORTANT TYPO IN HEADER - says SDSSJ0833+2040 - but object is actually SDSSJ0833+1204)

UT 09:39 - Object too faint for shift image centroid, blob at 1005,1030, will shift 19pixels = +2.24"

UT 09:51 - Executing science script

UT 09:53 - Had to manually start Flexure compensation

UT 09:55 - Seeing 0.73" Guider., 0.77" DIMM
UT 10:07 - +0,30" to move back to position 1 - doing a 3rd exposure

UT 10:20 - moving to Telluric std

UT 10:31 - spectra of Telluric

UT 10:36 - Completed Science 2 (Bulgless AGN)

09:34 guider .9 wfsc .66
  • right_guiderimage000179.fits 0.9

  • gpfocus -1
  • right_guiderimage000189.fits 1.0

  • gpfocus 0
  • right_guiderimage000194.fits 0.9

  • pgfocus +1
  • right_guiderimage000202.fits .85

  • gpfocus +2
  • right_guiderimage000215.fits .8

  • gpfocus +3
  • right_guiderimage000225.fits 0.8

  • gpfocus 0
  • right_guiderimage000236.fits 0.9

  • gpfocus +3
  • right_guiderimage000249.fits 0.75

  • gpfocus 0
  • right_guiderimage000267.fits .8

  • gpfocus 5
  • right_guiderimage000273.fits .7

  • gpfocus 0
  • right_guiderimage000292.fits 0.9

  • pgfocus 5
  • right_guiderimage000304.fits .7

  • gpfocus 0
  • right_guiderimage000315.fits .9

  • gpfocus 5
  • right_guiderimage000323.fits .75
  • lbto@obs4:4 % isisCmd --mods2 m2.ie 'gpfocus'
  • DONE: GPFOCUS AGWFOCUS=5.000 AGWFS=56.332 AGWFS0=39.504 AGWYS=29.624

  • gpfocus 0
  • lbto@obs4:5 % isisCmd --mods2 m2.ie 'gpfocus'
  • DONE: GPFOCUS AGWFOCUS=0.000 AGWFS=51.332 AGWFS0=39.504 AGWYS=34.624

09:55 Done

10:01 StopWFSing

10:01 collimate_multiple, /SEND_COW, /SEND_ZERN

10:12 Add 500 Z5

10:14 Add 500 Z9

10:15 New star. Telloric for Barry. Central star is too bright for guiding. Off-axis star is too far off-axis for MODS2 guide probe. 10 minutes so so will stop MODS2 stuff while Barry takes his data.

10:40 Preset to Stone_E

10:30 start transform_collect

10:45 send preset to left only. Get bad guide image on MODS2
  • right_guiderimage000579.fits
  • right_guiderimage000580.fits

11:46 Left primary went down.
  • Right guide star looked bad. Took 4 iterations to collimate.
  • right_wfscimage 127-130.
  • May not be related to SX primary. Then next on-axis collimation took two iterations

13:12 finish transform_collect with 27 points

13:12 Stop WFSing, start collimate_multiple with dynamic recon.

13:15:clear active optics on prim and sec. Start collimate_multiple on wfscimage 25

TTM log in /home/LBTO/data/20141128/ttm/dice5_luci.log

13:30 Geno close

-- %USERSIG{BarryRothberg - 2014-11-28}%

Comments

Topic revision: r3 - 28 Nov 2014, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback